OpenStudy (anonymous):

Find the exact value of the trigonometric function. cot(- 5pi/4) A. -1 B. - sq root of 2 C. 1 D. - sq root of 2/2

jimthompson5910 (jim_thompson5910):

hint: cot(x) = 1/tan(x) so cot(-5pi/4) = 1/tan(-5pi/4)

jimthompson5910 (jim_thompson5910):

since tan(x) is an odd function, you can also say cot(-5pi/4) = 1/tan(-5pi/4) cot(-5pi/4) = 1/(-tan(5pi/4)) cot(-5pi/4) = -1/tan(5pi/4)